The Linux Keyboard HOWTO

Andries Brouwer, aeb@cwi.nl v2.1, 8 November 1995

This note contains some information about the Linux keyboard and console, and the use of non-ASCII characters.

Contents

1 Useful programs 1

2 Keyboard generalities 1

3 Console generalities 2

4 Resetting your terminal 2 4.1 Keyboard hardware reset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

5 Delete and Backspace 3

5.1 How to tell Unix what character you want to use to delete the last typed character . . . . . . 3 5.1.1 ‘Getty used to do the right thing with DEL and BS but is broken now?’ . . . . . . . . 3 5.1.2 ‘Login behaves differently at the first and second login attempts?’ . . . . . . . . . . . 3

5.2 How to tell Linux what code to generate when a key is pressed . . . . . . . . . . . . . . . . . 3

5.2.1 ‘Why doesn’t the Backspace key generate BackSpace by default?’ . . . . . . . . . . . . 4 5.3 How to tell X to interchange Delete and Backspace . . . . . . . . . . . . . . . . . . . . . . . . 4 5.4 How to tell emacs what to do when it receives a Delete or Backspace . . . . . . . . . . . . . . 4 5.5 How to tell emacs to interchange Delete and Backspace . . . . . . . . . . . . . . . . . . . . . 5 5.6 How to tell kermit to interchange Delete and Backspace . . . . . . . . . . . . . . . . . . . . . 5 5.7 How to tell xterm about your favourite tty modes . . . . . . . . . . . . . . . . . . . . . . . . . 5 5.8 How to tell xmosaic that the Backspace key generates a DEL . . . . . . . . . . . . . . . . . . 5 5.9 What about termcap and terminfo? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

6 The console character sets 6

7 Console switching 7 7.1 Changing the number of Virtual Consoles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 8 Ctrl-Alt-Del and other special key combinations 8

8.1 Ctrl-Alt-Del (Boot) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 8.2 Other combinations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 8.3 X Combinations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 8.4 Dosemu Combinations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 8.5 Composing symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

9 Various properties of the console 10 10 How to get out of raw mode 10 11 The TERM variable 11

11.1 Terminfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 12 How to make other programs work with non-ASCII chars 11 13 What precisely does XFree86-2.1 do when it initializes its keymap? 13 14 Unusual keys and keyboards 13 15 Examples of use of loadkeys and xmodmap 14

15.1 ‘I can use only one finger to type with’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 16 Changing the video mode 15

16.1 Instructions for the use of resizecons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 17 Changing the keyboard repeat rate 16 18 Screensaving 16 19 Some properties of the VT100 - application key mode 16 20 Hardware incompatibility 17

1 Useful programs

The following packages contain keyboard or console related programs.

kbd-0.90.tar.gz contains loadkeys, dumpkeys, showkey, setmetamode, setleds, setfont, showfont,
mapscrn, kbd mode, chvt, resizecons, disalloc, getkeycodes, setkeycodes.
util-linux-2.5 contains setterm, kbdrate.
shellutils-1.8 (or stty.tar.gz) contains stty.
open-1.1.tgz contains open. (See also dynamic-vc-1.1.tar.gz.)
SVGATextMode-0.9.tar.gz contains SVGATextMode.

The X distribution contains xmodmap, xset, kbd mode. (See also X386keybd(1).)

2 Keyboard generalities

You press a key, and the keyboard controller sends scancodes to the kernel keyboard driver. Some keyboards can be programmed, but usually the scancodes corresponding to your keys are fixed. The kernel keyboard driver just transmits whatever it receives to the application program when it is in scancode mode, like when X is running. Otherwise, it parses the stream of scancodes into keycodes, corresponding to key press or key release events. (A single key press can generate up to 6 scancodes.) These keycodes are transmitted to the application program when it is in keycode mode (as used, for example, by showkey). Otherwise, these keycodes are looked up in the keymap, and the character or string found there is transmitted to the application, or the action described there is performed. (For example, if one presses and releases the a key, then the keyboard produces scancodes 0x1e and 0x9e, this is converted to keycodes 30 and 158, and then transmitted as 0141, the ASCII or latin-1 code for ’a’; if one presses and releases Delete, then the keyboard produces scancodes 0xe0 0x53 0xe0 0xd3, these are converted to keycodes 111 and 239, and then transmitted as the 4-symbol sequence ESC [ 3 ~, all assuming a US keyboard and a default keymap. An example of a key combination to which an action is assigned is Ctrl-Alt-Del.)

The translation between unusual scancodes and keycodes can be set using the utility setkeycodes - only very few people will need it. The translation between keycodes and characters or strings or actions, that is, the keymap, is set using the utility loadkeys. For details, see getkeycodes(8), setkeycodes(8), dumpkeys(1), loadkeys(1).

Where it says ‘transmitted to the application’ in the above description, this really means ‘transmitted to the terminal driver’. That is, further processing is just like that of text that comes in over a serial line. The details of this processing are set by the program stty.

3 Console generalities

Conversely, when you output something to the console, it first undergoes the standard tty processing, and then is fed to the console driver. The console driver emulates a VT100, and parses the input in order to recognize VT100 escape sequences (for cursor movement, clear screen, etc.). The characters that are not part of an escape sequence are transformed using a mapping table, and the transformed bytes are written in video memory, where they cause the display of character shapes found in the video card’s character ROM.

One can load one’s own fonts into character ROM using setfont, and load the user mapping table using mapscrn. More details will be given below.

4 Resetting your terminal

There is garbage on the screen, or all your keystrokes are echoed as line drawing characters. What to do?

Many programs will redraw the screen when ^L is typed. This might help when there is some modem noise or broadcast message on your screen. The command clear will clear the screen.

The command reset will reset the console driver. This helps when the screen is full of funny graphic characters, and also if it is reduced to the bottom line. If you don’t have this command, or if it does something else, make your own by putting the following two lines in an executable file reset in your PATH:

     #!/bin/sh
     echo -e \\033c

that is, you want to send the two characters ESC c to the console. If you loaded some strange font, and want to return to the default,

     % setfont

will do (provided you stored the default font in the default place). On old terminals output involving tabs may require a delay, and you have to say

     % stty tab3

(see stty(1)). You can change the video mode using resizecons or SVGATextMode. This usually settles the output side. On the input side there are many things that might be wrong. If X or DOOM or some other program using raw mode crashed, your keyboard may still be in raw (or mediumraw) mode, and it is difficult to give commands. (See ”How to get out of raw mode” below.)

4.1 Keyboard hardware reset

Things may be wrong on a lower level than Linux knows about. There are at least two distinct lower levels (keyboard and keyboard controller) where one can give the command ”keyboard disable” to the keyboard hardware. Keyboards can often be programmed to use one out of three different sets of scancodes.

However, I do not know of cases where this turned out to be a problem.

Some keyboards have a remapping capability built in. Stormy Henderson (stormy@Ghost.Net) writes: ‘If it’s your keyboard accidently being reprogrammed, you can (on a Gateway AnyKey keyboard) press controlalt-suspend macro to reset the keys to normal.’

5 Delete and Backspace

5.1 How to tell Unix what character you want to use to delete the last typed character

     % stty erase ^?

If the character is erased, but in a funny way, then something is wrong with your tty settings. If echoprt is set, then erased characters are enclosed between \ and /. If echoe is not set, then the erase char is echoed (which is reasonable when it is a printing character, like #). Most people will want stty echoe -echoprt. Saying stty sane will do this and more. Saying stty -a shows your current settings. How come this is not right by default? It is, if you use the right getty.

Note that many programs (like bash, emacs etc.) have their own keybindings (defined in ~/.inputrc, ~/.emacs etc.) and are unaffected by the setting of the erase character.

5.1.1 ‘Getty used to do the right thing with DEL and BS but is broken now?’

Earlier, the console driver would do BS Space BS (\010\040\010) when it got a DEL (\177). Nowadays, DEL’s are ignored (as they should be, since the driver emulates a vt100). Get a better getty, i.e., one that does not output DEL.

5.1.2 ‘Login behaves differently at the first and second login attempts?’

At the first attempt, you are talking to getty. At the second attempt, you are talking to login, a different program.

5.2 How to tell Linux what code to generate when a key is pressed

On the console, or, more precisely, when not in (MEDIUM)RAW mode, use

     % loadkeys mykeys.map

and under X use

     % xmodmap mykeys.xmap

Note that (since XFree86-2.1) X reads the Linux settings of the keymaps when initialising the X keymap. Although the two systems are not 100% compatible, this should mean that in many cases the use of xmodmap has become superfluous.

If, for example, you want your Backspace key to generate BackSpace (^H), instead of the default Delete, then 5. Delete and Backspace 6

% loadkeys
keycode 14 = BackSpace
%

will do.

5.2.1 ‘Why doesn’t the Backspace key generate BackSpace by default?’

(i) Because the VT100 had a Delete key above the Enter key.

(ii) Because Linus decided so.

5.3 How to tell X to interchange Delete and Backspace

% xmodmap -e "keysym BackSpace = Delete" -e "keysym Delete = BackSpace"

Or, if you just want the Backspace key to generate a BackSpace:

% xmodmap -e "keycode 22 = BackSpace"

Or, if you just want the Delete key to generate a Delete:

% xmodmap -e "keycode 107 = Delete"

(but usually this is the default binding already).

5.4 How to tell emacs what to do when it receives a Delete or Backspace

Put in your .emacs file lines like

(global-set-key "\?" 'delete-backward-char)
(global-set-key "\C-h" 'help-command)

Of course you can bind other commands to other keys in the same way.

5.5 How to tell emacs to interchange Delete and Backspace

Put in your .emacs file lines

(setq keyboard-translate-table (make-string 128 0))
(let ((i 0))
(while (< i 128)
    (aset keyboard-translate-table i i)
    (setq i (1+ i))))
(aset keyboard-translate-table ?\b ?\^?)
(aset keyboard-translate-table ?\^? ?\b)

5.6 How to tell kermit to interchange Delete and Backspace

Put in your .kermrc file the lines

set key \127 \8
set key \8 \127

5.7 How to tell xterm about your favourite tty modes

Normally xterm will inherit the tty modes from its invoker. Under xdm, the default erase and kill characters are # and @, as in good old Unix Version 6. If you don’t like that, you might put something like

XTerm*ttymodes: erase ^? kill ^U intr ^C quit ^\ eof ^D susp ^Z start ^Q stop ^S eol ^@
in /usr/lib/X11/app-defaults/XTerm or in $HOME/.Xresources, assuming that you have a line
xrdb $HOME/.Xresources
in your $HOME/.xinitrc.

5.8 How to tell xmosaic that the Backspace key generates a DEL

Putting

*XmText.translations: #override\n\
     <Key>osfDelete: delete-previous-character()
*XmTextField.translations: #override\n\
      <Key>osfDelete: delete-previous-character()
in your $HOME/.Xresources helps.

The netscape FAQ, however, says:

Why doesn't my Backspace key work in text fields?
By default, Linux and XFree86 come with the Backspace and Delete keys
misconfigured. All Motif programs (including, of course, Netscape
Navigator) will malfunction in the same way.
The Motif spec says that Backspace is supposed to delete the previous
character and Delete is supposed to delete the following character.
Linux and XFree86 come configured with both the Backspace and Delete
keys generating Delete.
You can fix this by using any one of the xmodmap, xkeycaps, or
loadkeys programs to make the key in question generate the BackSpace

        keysym instead of Delete.
        You can also fix it by having a .motifbind file; see the man page
        for VirtualBindings(3).
        Note: Don't use the *XmText.translations or *XmTextField.translations
        resources to attempt to fix this problem. If you do, you will blow away
        Netscape Navigator's other text-field key bindings.

5.9 What about termcap and terminfo?

When people have problems with backspace, they tend to look at their termcap (or terminfo) entry for the terminal, and indeed, there does exist a kb (or kbs) capability describing the code generated by the Backspace key. However, not many programs use it, so unless you are having problems with one particular program only, probably the fault is elsewhere. Of course it is a good idea anyway to correct your termcap (terminfo) entry. See also below under ”The TERM variable”.

6 The console character sets

The kernel knows about 4 translations of bytes into console-screen symbols. The four tables are: a) Latin1 -> PC, b) VT100 graphics -> PC, c) PC -> PC, d) user-defined.

There are two character sets, called G0 and G1, and one of them is the current character set. (Initially G0.) Typing ^N causes G1 to become current, ^O causes G0 to become current.

These variables G0 and G1 point at a translation table, and can be changed by the user. Initially they point at tables a) and b), respectively. The sequences ESC ( B and ESC ( 0 and ESC ( U and ESC ( K cause G0 to point at translation table a), b), c) and d), respectively. The sequences ESC ) B and ESC ) 0 and ESC ) U and ESC ) K cause G1 to point at translation table a), b), c) and d), respectively.

The sequence ESC c causes a terminal reset, which is what you want if the screen is all garbled. The oftadvised echo ^V^O will only make G0 current, but there is no guarantee that G0 points at table a). In some distributions there is a program reset(1) that just does echo ^[c. If your termcap entry for the console is correct (and has an entry :rs=\Ec:), then also setterm -reset will work.

The user-defined mapping table can be set using mapscrn(8). The result of the mapping is that if a symbol c is printed, the symbol s = map[c] is sent to the video memory. The bitmap that corresponds to s is found in the character ROM, and can be changed using setfont(8).

7 Console switching

By default, console switching is done using Alt-Fn or Ctrl-Alt-Fn. Under X (or recent versions of dosemu), only Ctrl-Alt-Fn works.

XFree86 1.3 does not know that Alt is down when you switch to the X window. Thus, you cannot switch immediately to some other VT again but have to release Alt first. In the other direction this should work: the kernel always keeps track of the up/down status of all keys. (As far as possible: on some keyboards some keys do not emit a scancode when pressed (e.g.: the PFn keys of a FOCUS 9000) or released (e.g.: the Pause key of many keyboards).)

XFree86 1.3 saves the fonts loaded in the character ROMs when started, and restores it on a console switch. Thus, the result of setfont on a VT is wiped out when you go to X and back. Using setfont under X will lead to funny results.

One can change under program control using the chvt command.

7.1 Changing the number of Virtual Consoles

This question still comes up from time to time, but the answer is: you already have enough of them. Since kernel version 1.1.54, there are between 1 and 63 virtual consoles. A new one is created as soon as it is opened. It is removed by the utility disalloc (but it can be removed only when no processes are associated to it anymore).

For older kernels, change the line

     #define NR_CONSOLES     8

in include/linux/tty.h (don’t increase this number beyond 63), and recompile the kernel. You might have to create the devices with MAKEDEV or mknod ttyN c 4 N where N denotes the tty number. If you want the new VCs to run getty, add lines in /etc/inittab.

When the consoles are allocated dynamically, it is usually easiest to have only one or two running getty. More are opened by open -l -s bash. Unused consoles (without associated processes) are disallocated

using disalloc.
If you have spawn login from kbd-0.90.tar.gz and you put
             loadkeys << EOF
             alt keycode 103 = Spawn_Console
             EOF
             spawn_login &

in /etc/rc.local, then typing Alt-UpArrow will create a fresh VC running login (and switch to it). See

also open-1.3.tgz.

You can only login as ”root” on terminals listed in /etc/securetty.

8 Ctrl-Alt-Del and other special key combinations

8.1 Ctrl-Alt-Del (Boot)

If you press Ctrl-Alt-Del (or whatever key was assigned the keysym Boot by loadkeys) then either the machine reboots immediately (without sync), or init is sent a SIGINT. The former behaviour is the default. The default can be changed by root, using the system call reboot(), see ctrlaltdel(8). Some init’s change the default. What happens when init gets SIGINT depends on the version of init used - often it will be determined by the pf entry in /etc/inittab (which means that you can run an arbitrary program in this case). In the current kernel Ctrl-AltGr-Del is no longer by default assigned to Boot.

8.2 Other combinations

Name            Default binding
-------------------------------
Show_Memory     Shift-Scrollock
Show_Registers  AltGr-ScrollLock
Show_State      Ctrl-ScrollLock
Console_n       Alt-Fn and Ctrl-Alt-Fn  (1 <= n <= 12)
Console_{n+12}  AltGr-Fn                (1 <= n <= 12)
Incr_Console    Alt-RightArrow
Decr_Console    Alt-LeftArrow
Last_Console    Alt[Gr]-PrintScreen
Scroll_Backward Shift-PageUp
Scroll_Forward  Shift-PageDown
Caps_On                                 (CapsLock is a toggle; this key sets)
Compose         Ctrl-.

8.3 X Combinations

Ctrl-Alt-Fn     Switch to VT n
Ctrl-Alt-KP+    Next mode
Ctrl-Alt-KP-    Previous mode
Ctrl-Alt-Backspace      Kill X

On some motherboards, Ctrl-Alt-KP- and Ctrl-Alt-KP+ will be equivalent to pressing the Turbo button. That is, both will produce the scancodes 1d 38 4a ca b8 9d and 1d 38 4e ce b8 9d, and both will switch between Turbo (>= 25MHz) and non-Turbo (8 or 12 MHz). (Often these key combinations only function this way when enabled by jumpers on the motherboard.)

Perry F Nguyen (pfnguyen@netcom22.netcom.com) writes: AMI BIOS has a feature that locks up the keyboard and flashes the LED’s if the Ctrl-Alt-Backspace combination is pressed while a BIOS password is enabled, until the CMOS/BIOS password is typed in.

8.4 Dosemu Combinations

Ctrl-Alt-Fn     Switch to VT n (from version 0.50; earlier Alt-Fn)
Ctrl-Alt-PgDn   Kill dosemu (when in RAW keyboard mode)
(and many other combinations - see the dosemu documentation)

8.5 Composing symbols

One symbol may be constructed using several keystrokes.

• LeftAlt-press, followed by a decimal number typed on the keypad, followed by LeftAlt-release, yields the symbol with code given by this number. (In Unicode mode this same mechanism, but then with 4 hexadecimal digits, may be used to define a Unicode symbol.)

• A dead diacritic followed by a symbol, yields that symbol adorned with that diacritic. If the combination is undefined, both keys are taken separately. Which keys are dead diacritics is user-settable; none is by default. Five dead diacritics can be defined (using loadkeys(1)): dead grave, dead acute, dead circumflex, dead tilde, dead diaeresis. Precisely what this adorning means is also user-settable: dead-diacritic, symbol is equivalent to Compose + diacritic + symbol.

• Compose followed by two symbols yields a combination symbol. These combinations are user-settable. Today there are 68 combinations defined by default; you can see them by saying ”dumpkeys | grep compose”.

• Then there are ‘Sticky’ modifier keys (since 1.3.33). For example, one can type ^C as SControl, C and Ctrl-Alt-BackSpace as SControl, SAlt, BackSpace.

Note that there are at least three such composition mechanisms:

1. The Linux keyboard driver mechanism, used in conjunction with loadkeys.

2. The X mechanism - see X386keybd(1), later XFree86kbd(1). Under X11R6: edit

     /usr/X11R6/lib/X11/locale/iso8859-1/Compose.

3. The emacs mechanism obtained by loading ”iso-insert.el”.

For X the order of the two symbols is arbitrary: both Compose-,-c and Compose-c-, yield a c-cedilla; for Linux and emacs only the former sequence works by default. For X the list of compose combinations is fixed. Linux and emacs are flexible. The three default lists are somewhat similar, but the details are different.

9 Various properties of the console

See loadkeys(1), setleds(1), setmetamode(1) for the codes generated by the various keys and the setting of leds when not under X. Under X, see xmodmap(1).

See setterm(1), kbdrate(8) for properties such as foreground and background colors, screen blanking, character repeat rate when not under X. Under X, see xset(1), also for key click and bell volume.

The file /etc/termcap defines the escape sequences used by many programs addressing the console (or any other terminal). A more modern version is found in /usr/lib/terminfo. (See terminfo(5). Terminfo files are compiled by the terminfo compiler /usr/lib/terminfo/tic, see tic(1).)

(On my machine) /dev/console is a symbolic link to /dev/tty0, and the kernel regards /dev/tty0 as a synonym for the current VT. XFree86 1.3 changes the owner of /dev/tty0, but does not reset this after finishing. Thus, dumpkeys might fail because someone else owns /dev/tty0; in such a case you might run X first.

10 How to get out of raw mode

If some program using K RAW keyboard mode exits without restoring the keyboard mode to K XLATE, then it is difficult to do anything - not even Ctrl-Alt-Del works. However, it is sometimes possible to avoid hitting the reset button. (And desirable as well: your users may get angry if you kill their Hack game by rebooting; you might also damage your file system.) Easy solutions involve logging in from another terminal or another machine and doing kbd mode -a. The procedure below assumes that no X is running, that the display is in text mode, and that you are at your bash prompt, that you are using a US keyboard layout, and that your interrupt character is Ctrl-C.

Step 1. Start X. As follows: press 2 (and don’t release), press F12 (and don’t release) and immediately afterwards press = . This starts X. (Explanation: if a key press produces keycode K, then the key release produces keycode K+128. Probably your shell does not like these high characters, so we avoid generating them by not releasing any key. However, we have to be quick, otherwise key repeat starts. The digit 2 produces a Ctrl-C that discards previous junk, the F12 produces an X and the = a Return.) Probably your screen will be grey now, since no .xinitrc was specified. However, Ctrl-Alt-Fn will work and you can go to another VT. (Ctrl-Alt-Backspace also works, but that exits X, and gets you back into the previous state, which is not what you want.)

Step 2. Setup to change the keyboard mode. (For example, by sleep 5; kbd mode -a.)

Step 3. Leave X again. Alt-Fx (often Alt-F7) brings you back to X, and then Ctrl-Alt-Backspace exits X. Within 5 seconds your keyboard will be usable again.

If you want to prepare for the occasion, then make \215A\301 (3 symbols) an alias for kbd mode -a. Now just hitting = F7 = (3 symbols) will return you to sanity.

11 The TERM variable

Many programs use the TERM variable and the database /etc/termcap or /usr/lib/terminfo/* to decide which strings to send for clear screen, move cursor, etc., and sometimes also to decide which string is sent by the users backspace key, function keys etc. This value is first set by the kernel (for the console). Usually, this variable is re-set by getty, using /etc/ttytype or the argument specified in /etc/inittab. Sometimes, it

is also set in /etc/profile.

Older systems use TERM=console or TERM=con80x25. Newer systems (with ncurses 1.8.6) use the more specific TERM=linux or TERM=linux-80x25. However, old versions of setterm test for TERM=con* and hence fail to work with TERM=linux.

Since kernel version 1.3.2, the kernel default for the console is TERM=linux.

If you have a termcap without entry for linux, add the word linux to the entry for the console:

             console|con80x25|linux:\
and make /usr/lib/terminfo/l/linux a copy of or symbolic link to /usr/lib/terminfo/c/console.

11.1 Terminfo

The terminfo entry for the linux console from ncurses 1.8.6 misses the entry kich1=\E[2~, needed by some programs. Edit the file and tic it.

12 How to make other programs work with non-ASCII chars

In the bad old days this used to be quite a hassle. Every separate program had to be convinced individually to leave your bits alone. Not that all is easy now, but recently a lot of gnu utilities have learned to react to LC CTYPE=iso 8859 1 or LC CTYPE=iso-8859-1. Try this first, and if it doesn’t help look at the hints below.

First of all, the 8-th bit should survive the kernel input processing, so make sure to have stty cs8 -istrip

-parenb set.

A. For emacs, put lines

             (standard-display-european t)
             (set-input-mode nil nil 1)
             (require 'iso-syntax)

and perhaps also

             (load-file "iso-insert.el")
             (define-key global-map [?\C-.] 8859-1-map)

into your $HOME/.emacs. (The latter line will not work under xterm, if you use emacs -nw, but in that case you can put

             XTerm*VT100.Translations:       #override\n\
             Ctrl <KeyPress> . : string("\0308")
in your .Xresources.)

B. For less, put LESSCHARSET=latin1 in the environment.

C. For ls, give the option -N. (Probably you want to make an alias.)

D. For bash (version 1.13.*), put

             set meta-flag on
             set convert-meta off

and, according to the Danish HOWTO,

             set output-meta on
into your $HOME/.inputrc.

E. For tcsh, use

             setenv LANG     US_en
             setenv LC_CTYPE iso_8859_1

If you have nls on your system, then the corresponding routines are used. Otherwise tcsh will assume iso 8859 1, regardless of the values given to LANG and LC CTYPE. See the section NATIVE LANGUAGE SYSTEM in tcsh(1). (The Danish HOWTO says: setenv LC CTYPE ISO-8859-1; stty pass8)

F. For flex, give the option -8 if the parser it generates must be able to handle 8-bit input. (Of course it must.)

G. For elm, set displaycharset to ISO-8859-1. (Danish HOWTO: LANG=C and LC CTYPE=ISO-8859-1)

H. For programs using curses (such as lynx) David Sibley reports: The regular curses package uses the highorder bit for reverse video mode (see flag STANDOUT defined in /usr/include/curses.h). However, ncurses seems to be 8-bit clean and does display iso-latin-8859-1 correctly.

I. For programs using groff (such as man), make sure to use -Tlatin1 instead of -Tascii. Old versions of the program man also use col, and the next point also applies.

J. For col, make sure 1) that it is fixed so as to do setlocale(LC CTYPE,""); and 2) to put

LC CTYPE=ISO-8859-1 in the environment.

K. For rlogin, use option -8.

L. For joe, sunsite.unc.edu:/pub/Linux/apps/editors/joe-1.0.8-linux.tar.gz is said to work after

editing the configuration file. Someone else said: joe: Put the -asis option in /isr/lib/joerc in the first column.

M. For LaTeX: \documentstyle

isolatin

{article}. For LaTeX2e: \documentclass{article}\usepackage{isolatin} where isolatin.sty is
available from ftp://ftp.vlsivie.tuwien.ac.at/pub/8bit.

A nice discussion on the topic of ISO-8859-1 and how to manage 8-bit characters is contained in the file grasp.insa-lyon.fr:/pub/faq/fr/accents (in French). Another fine discussion (in English) can

be found in rtfm.mit.edu:pub/usenet-by-group/comp.answers/character-sets/iso-8859-1-faq. And
another(?), in ftp.vlsivie.tuwien.ac.at:/pub/8bit/FAQ-ISO-8859-1.

13 What precisely does XFree86-2.1 do when it initializes its keymap?

Since version 2.1, XFree86 will initialize its keymap from the Linux keymap, as far as possible. However, Linux had 16 entries per key (one for each combination of the Shift, AltGr, Ctrl, Alt modifiers) and presently has 256 entries per key, while X has 4 entries per key (one for each combination of Shift, Mod), so some information is necessarily lost.

First X reads the Xconfig file, where definitions of the LeftAlt, RightAlt, RightCtl, ScrollLock keys as Meta, ModeShift, Compose, ModeLock or ScrollLock might be found - see X386keybd(1), later XFree86kbd(1).

For Mod the LeftAlt key is taken, unless RightCtl was defined as ModeShift or ModeLock, in which case RightCtl is taken, or RightAlt was so defined, in which case RightAlt is taken. This determines how the 4 XFree86 meanings of a key are selected from the 16 Linux meanings. Note that Linux today does not distinguish by default between the two Ctrl keys or between the two Shift keys. X does distinguish.

Now the kernel keymap is read and the usually obvious corresponding X bindings are made. The bindings for the ”action keys” Show Memory, Show State, Show Registers, Last Console, Console n, Scroll Backward, Scroll Forward, Caps On and Boot are ignored, as are the dead diacriticals, and the locks (except for ShiftLock), and the ”ASCII-x” keys.

Next, the definitions in the Xconfig file are used. (Thus, a definition of Compose in Xconfig will override its value as found in the Linux keymap.)

What happens to the strings associated with the function keys? Nothing, X does not have such a concept. (But it is possible to define strings for function keys in xterm - note however that the window manager gets the keys first.)

I don’t know how to convince xterm that it should use the X keymap when Alt is pressed; it seems just to look at its resource eightBitInput, and depending on whether that is true or false either set the high order bit of the character, or generate an additional Escape character (just like setmetamode(1) does for the console).

14 Unusual keys and keyboards

The two keys PrintScrn/SysRq and Pause/Break are special in that they have two keycodes: the former has keycode 84 when Alt is pressed simultaneously, and keycode 99 otherwise; the latter has keycode 101 when Ctrl is pressed simultaneously, and keycode 119 otherwise. (Thus, it makes no sense to bind functions to Alt keycode 99 or Ctrl keycode 119.)

If you have strange keys, that do not generate any code under Linux (or generate messages like ”unrecognized scancode”), and your kernel is 1.1.63 or later, then you can use setkeycodes(1) to tell the kernel about them. They won’t work under X, however. Once they have gotten a keycode from setkeycodes, they can be assigned a function by loadkeys.

15 Examples of use of loadkeys and xmodmap

Switching Caps Lock and Control on the keyboard (assuming you use keymaps 0-15; check with dumpkeys

| head -1)
       % loadkeys
       keymaps 0-15
       keycode 58 = Control
       keycode 29 = Caps_Lock
       %

Switching them under X only:

       % xmodmap .xmodmaprc

where .xmodmaprc contains lines

       remove Lock = Caps_Lock
       remove Control =  Control_L
       keysym  Control_L  =  Caps_Lock
       keysym  Caps_Lock  = Control_L
       add Lock = Caps_Lock
       add Control = Control_L

What is this about the key numbering? Backspace is 14 under Linux, 22 under X? Well, the numbering can best be regarded as arbitrary; the Linux number of a key can be found using showkey(1), and the X number using xev(1). Often the X number will be 8 more than the Linux number.

15.1 ‘I can use only one finger to type with’

Can the Shift, Ctrl and Alt keys be made to behave as toggles?” Yes, after saying

     % loadkeys
     keycode 29 = Control_Lock
     keycode 42 = Shift_Lock
     keycode 56 = Alt_Lock
     %

the left Control, Shift and Alt keys will act as toggles. The numbers involved are revealed by showkey (and usually are 29, 97, 42, 54, 56, 100 for left and right control, shift and alt, respectively), and the functions are Control Lock, Shift Lock, Alt Lock, ALtGr Lock.

What about ‘sticky’ modifier keys? Well, since their introduction in the kernel (1.3.33) there has not been a new release of the kbd package yet, so they have to be invoked using their hexadecimal codes. For example,

     % loadkeys
     keymaps 0-15

     keycode 54 = 0x0c00
     keycode 97 = 0x0c02
     keycode 100 = 0x0c03
     %

will make the right Shift, Ctrl, Alt sticky versions of the left ones.

16 Changing the video mode

As far as I know there are 6 ways to change resolution:

1. At compile time: change the line

             SVGA_MODE=      -DSVGA_MODE=NORMAL_VGA
in /usr/src/linux/Makefile.

1A. After compilation: use rdev -v - a terrible hack, but it exists.

2. At boot time: put vga=ask in the lilo config file, and lilo will ask you what video mode you want. Once you know, put vga=mypreference.

3. At run time: A. Use the resizecons command. (This is a very primitive wrapper around the VT RESIZE ioctl.) B. Use the SVGATextMode command. (This is a less primitive wrapper around the VT RESIZE ioctl.)

4. Not ”on the console”: Under dosemu, or with svgalib etc. you can change the hardware video mode without the console driver being aware of it. Sometimes this is useful in getting resizecons or SVGATextMode set up: use dosemu and some DOS program to get into the desired videomode, dump (say from another VT) the contents of all video hardware registers, and use that in the initialization that resizecons and SVGATextMode require. In some cases where the video mode has gotten into some unusable state, starting dosemu, relying on the BIOS to set up the video mode, and then killing dosemu (with kill -9), is the easiest way to get into shape again.

16.1 Instructions for the use of resizecons

Get svgalib and compile the program restoretextmode. Boot up your machine in all possible video modes (using vga=ask in the lilo config file), and write the video hardware register contents to files CxR (C=cols, R=rows), e.g., 80x25, 132x44, etc. Put these files in /usr/lib/kbd/videomodes. Now resizecons 132x44 will change videomode for you (and send SIGWINCH to all processes that need to know about this, and load another font if necessary).

At present, resizecons only succeeds when there is memory enough for both the old and the new consoles at the same time.

17 Changing the keyboard repeat rate

At startup, the Linux kernel sets the repeat rate to its maximal value. For most keyboards this is reasonable, but for some it means that you can hardly touch a key without getting three copies of the corresponding symbol. Use the program kbdrate(8) to change the repeat rate, or, if that doesn’t help, edit or remove the section

     ! set the keyboard repeat rate to the max
         mov     ax,#0x0305
         xor     bx,bx           ! clear bx
         int     0x16
of /usr/src/linux/

arch/i386/

boot/setup.S.

18 Screensaving

setterm -blank nn will tell the console driver to blank the screen after nn minutes of inactivity. (With nn = 0, screensaving is turned off. In some old kernels this first took effect after the next keyboard interrupt.)

The s option of xset(1) will set the X screensaving parameters.

The video hardware powersaving modes can be enabled/disabled using the setvesablank program given in

the starting comment of /usr/src/linux/drivers/char/vesa blank.c.

19 Some properties of the VT100 - application key mode

: Sometimes my curser keys produce strange codes?

When the terminal is in application key mode the cursor keys produce Esc O x and otherwise Esc [ x where x is one of A,B,C,D. Certain programs put the terminal in application key mode; if you kill them with kill -9, or if they crash, then the mode will not be reset.

        % echo -e '\033c'

resets all properties of the current VC. Just changing the application key mode is done by

        % echo -e '\033[?1h'

(set) and

        % echo -e '\033[?1l'

(clear).

20 Hardware incompatibility

Several people have noticed that they lose typed characters when a floppy disk is active. It seems that this might be a problem with Uni-486WB motherboards. (Please mail me to confirm

yes, Ihavethesameproblem

, deny

no, nothingwrongwithmyUni − 486W B

, modify

MyXyzzymachinehasthesameproblem

.)

Some people experience sporadic lockups - sometimes associated to hard disk activity or other I/O.

ulf@rio70.bln.sni.de (Ulf Tietz) wrote: ‘I have had the same problems, when I had my motherboard tuned too fast. So I reset all the timings ( CLK, wait statements etc ) to more conventional values, and the problems are gone.’

bhogan@crl.com (Bill Hogan) wrote: ‘If you have an AMI BIOS, you might try setting the Gate A20 emulation parameter to ’chipset’ (if you have that option). Whenever I have had that parameter set to any of the other options on my machine (’fast’, ’both’, ’disabled’) I have had frequent keyboard lockups.’

——————————————————————–

Additions and corrections are welcome. Andries Brouwer - aeb@cwi.nl